# Online Cascade Learning for Efficient Inference over Streams

### v1

@@ 5.3 Adaptability to Larger Cascade
To further validate the adaptability of our online cascade learning system, we have also experimented with integrating a larger BERT-large model into the cascade. This augmentation aims to address more complex queries with greater precision while maintaining cost efficiency.
The results of this integration are encouraging, as shown in Appendix Figure 9 and 10. The larger cascade, particularly when trained on Llama 2 70B Chat annotations, demonstrates impressive performance across a variety of datasets, including IMDB, ISEAR, and FEVER. Notably, it achieves equivalent or even superior results compared to the standalone LLM in most cases, with the exception of the HateSpeech dataset where a larger cascade may overfit the task. This outcome underscores the cascade’s adaptability and its capacity for granular response tailoring to match the complexity levels of incoming queries.
The success of this larger cascade system signals the huge potential in our system’s evolution. It showcases our system’s inherent flexibility and scalability, enabling it to accommodate and efficiently utilize more powerful models as part of its architecture. Future investigations on online cascade systems may consider incorporating LLMs of different sizes and specialties, allowing for refined tailoring of responses to the specific characteristics of different queries.
@@ 6.1 Knowledge Distillation
Knowledge distillation, originally conceptualized by Hinton et al. (2015), emerged as a technique to transfer knowledge from a large, complex model (teacher) to a smaller, more efficient one (student), intending to retain performance while reducing computational costs. Notable advancements include the works of Sanh et al. (2019), who demonstrated the effectiveness of distilling the capabilities of BERT into smaller models, and Gu et al. (2023) who successfully applied distillation to LLMs, achieving comparable performance with significantly reduced model sizes. However, the significant difference in capabilities between the teacher and student model can lead to challenges, particularly when dealing with complex queries that require advanced reasoning or involve intricate subject matter (Cho & Hariharan, 2019; Zhang et al., 2022), thereby highlighting a performance gap that distillation alone cannot overcome.
@@ 6.3 Model Cascade
By orchestrating multiple models with varying complexities, model cascade has been widely adopted in both CV and NLP tasks to enhance system efficiency. An early theme in this field is the early exiting from a neural network’s intermediate layers (Liu et al., 2020; Xin et al., 2020; Schwartz et al., 2020). These works have later inspired the cascade of complete models (Li et al., 2020; Khalili et al., 2022). Among them, Varshney & Baral (2022) systematically examines the trade-off between accuracy and cost in a cascade of variants of BERT. Chen et al. (2023) further extends the cascade to cover multiple LLM APIs by incorporating a scoring function. Compared to the previous works, our work expands the model cascade by allowing small models to learn online and improve rather than having fixed capabilities. Moreover, by formulating the model confidence as part of the learning objective, our work eliminates the need to manually set confidence thresholds.
@@ 7 Conclusion
In this work, we address the challenge of managing streaming queries with LLMs in a cost-efficient way. We propose an online cascade learning framework that adapts to evolving queries by improving smaller models in the cascade through imitation of LLM behaviors. Our theoretical analysis provides a no-regret performance guarantee for the algorithm. Experimental results on several benchmarks confirmed the effectiveness of our approach, showing that it can achieve performance levels comparable to LLMs while significantly reducing inference costs, with potential savings of up to 90%.
@@ Appendix A Detailed Proofs and Theoretical Analysis
Given an online learning algorithm operating over time steps $1,...,T$, the regret $\gamma$ is defined as the difference between the total loss incurred by the algorithm and the loss of the best-fixed expert in hindsight. For online cascade learning, regret $\gamma$ is formally expressed as:
@@
We simplify by defining the expert policy as $\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi}J(\pi)$, leading to the following representation of regret:
@@ B.2 LLM Prompts
Sadness: Sadness is a feeling of sorrow, unhappiness, or despondency. Guilt: Guilt is a self-directed emotion that arises from a sense of wrongdoing or moral transgression.
@@
Fear: Fear is an emotion triggered by a perceived threat or danger. Disgust: Disgust is an aversive emotion linked to feelings of revulsion, repulsion, or strong distaste. It arises in response to things that are offensive or unpleasant.
@@ B.3 Experimental Configurations
The detailed hyperparameter settings for online cascade learning are listed in Table 2 and 3. Note that the “learning rate” in the table refers to the learning rates of the MLPs (in Section 3.2), not the models’. We used a consistent configuration for both online cascade learning and the distillation baselines by setting BERT-base’s batch size to 8, the learning rate to 0.00001, and the number of epochs to 5.
Table 2: Hyperparameter settings for online cascade learning experiments with GPT-3.5-Turbo as the LLM.

### v2

@@ 5.2 Case Analysis
Figure 9: Cost-accuracy trade-off curves on two input distributional shift scenarios, respectively using GPT-3.5 Turbo and Llama 2 70B Chat as the LLM with logistic regression and BERT-base in a cascade. “OCL” refers to online cascade learning, and “OEL” means online ensemble learning.
@@ 5.3 Adaptability to Larger Cascade
To further validate the adaptability of our online cascade learning system, we have also explored scaling up the cascade system by further integrating a BERT-large model (4 models in total). The results of this integration are encouraging, detailed in Appendix Figure 12 and 13. The expanded cascade system, particularly when trained on Llama 2 70B Chat annotations, demonstrates equivalent or even superior results compared to the standalone LLM in most cases, with the exception of the HateSpeech dataset, a simpler task setting where a larger cascade might complicate the deferral policy learning and thus degrading overall performance. Therefore, it is important to align cascade size with task complexity to optimize performance and avoid overfitting. A more detailed analysis is in Appendix C.2
The success of this larger cascade system signals the huge potential in our work’s extensions. It showcases our system’s inherent flexibility and scalability, enabling it to accommodate and efficiently utilize more powerful models as part of its architecture. Future investigations on online cascade systems may further scale up the system by incorporating LLMs of different sizes and specialties, allowing for refined tailoring of responses to the specific characteristics of different queries. However, there are also several factors worth noting when scaling up the cascade, as outlined in our analysis in Appendix C.1 and C.3.
5.4 Robustness against Distributional Shifts
Distribution Shift in Input Length
Longer inputs typically involve more complex semantics. For example, on IMDB, the average accuracy of GPT-3.5 Turbo is notably lower on longer movie reviews, as evidenced in Appendix Table 5. Therefore, to verify our method’s robustness against input distribution shifts, we rearrange the IMDB benchmark with length ascending order to simulate a distribution shift over the inputs’ semantic complexity. The results are shown in Figure 9 and Table 2. Despite the distribution shift in input length, online cascade learning demonstrates good performance with minimal accuracy drops across different cost budgets, regardless of the LLM adopted in the cascade.
Distribution Shift in Input Category
We further validate our approach with a distribution shift in input semantic categories, by filtering all the input samples regarding “Comedy” movies in the IMDB dataset (8,140 out of 25,000), and feeding them into the cascade as the last part of the input stream. This means that the system had not seen any comedy movie reviews in the first 2/3 of the inputs before processing comedy movie reviews in the last 1/3 of the inputs. As shown in Figure 9 and Table 2, our approach also performs reliably under the distribution shift in input categories, with a slight increase in average accuracy.
Based on the results, we conclude that our proposed approach can well utilize the advantages of online learning, quickly adapt to unseen inputs, and perform robustly against distribution shifts in the data streams.
GPT-3.5 Turbo Llama 2 70B
Without Any Distribution Shift 90.77% 90.97%
With Distribution Shift in Input Length 90.23% 90.64%
$\Rightarrow$ Difference -0.54% -0.33%
With Distribution Shift in Input Category 90.85% 91.46%
$\Rightarrow$ Difference +0.08% +0.49%
Table 2: Average accuracy of our approach across different cost budgets with distribution shifts in input length or input category, compared to the default setting.
@@ 6.1 Knowledge Distillation
Knowledge distillation, originally conceptualized by Hinton et al. (2015), emerged as a technique to transfer knowledge from a large, complex model (teacher) to a smaller, more efficient one (student), intending to retain performance while reducing computational costs. Notable advancements include the works of Sanh et al. (2019), who demonstrated the effectiveness of distilling the capabilities of BERT into smaller models, and Gu et al. (2023) who successfully applied distillation to LLMs, achieving comparable performance with significantly reduced model sizes. However, the significant difference in capabilities between the teacher and student model can lead to challenges, particularly when dealing with complex queries that require advanced reasoning or involve intricate subject matter (Cho & Hariharan, 2019; Zhang et al., 2022), thereby highlighting a performance gap that distillation alone cannot overcome (Rawat et al., 2021).
@@ 6.3 Model Cascade
By orchestrating multiple models with varying complexities, model cascade has been widely adopted in both CV and NLP tasks to enhance system efficiency. An early theme in this field is the early exiting from a neural network’s intermediate layers (Liu et al., 2020; Xin et al., 2020; Schwartz et al., 2020). These works have later inspired the cascade of complete models (Li et al., 2020; Khalili et al., 2022). Among them, Varshney & Baral (2022) systematically examines the trade-off between accuracy and cost in a cascade of variants of BERT. Chen et al. (2023) further extends the cascade to cover multiple LLM APIs by incorporating a scoring function. Compared to the previous works, our work expands the model cascade by allowing small models to learn online and improve rather than having fixed capabilities, similar to the recent idea of “neural caching” (Ramírez et al., 2023; Stogiannidis et al., 2023). Moreover, by formulating the model confidence as part of the learning objective, our work eliminates the need to set confidence thresholds manually.
@@ 7 Conclusion
In this work, we address the challenge of managing streaming queries with LLMs in a cost-efficient way. We propose an online cascade learning framework that adapts to evolving queries by improving smaller models in the cascade through imitation of LLM behaviors. Our theoretical analysis provides a no-regret performance guarantee for the algorithm. Extensive experiments confirmed the effectiveness of our approach, showing that it can achieve performance levels comparable to LLMs while significantly reducing inference costs, with potential savings of up to 90%.
@@
Acknowledgements
We thank the anonymous reviewers for their valuable comments and suggestions in enhancing the paper. This research was supported by the NSF under grant numbers CCF-1918651, 2008240, 2131294, and 2212557, ARO award #W911NF-21-1-0009, DARPA award #HR00112320018, NIH CTSA award No. UL1TR003167, and US DOT Tier-1 UTC CYBER-CARE grant #69A3552348332.
@@ Appendix A Detailed Proofs and Theoretical Analysis
Given an online learning algorithm operating over time steps $1,...,T$, the regret $\gamma$ is defined as the difference between the total loss incurred by the algorithm and the loss of the best fixed policy in hindsight. For online cascade learning, regret $\gamma$ is formally expressed as:
@@
We simplify by defining the best fixed policy as $\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi}J(\pi)$, leading to the following representation of regret:
@@ B.1 A Simple Prefill Experiment
Figure 10: A more detailed cost-performance trade-off plot with accuracy, F1-score, recall, and precision curves, using GPT-3.5 Turbo as the LLM.
Figure 11: A more detailed cost-performance trade-off plot with accuracy, F1-score, recall, and precision curves, using Llama 2 70B Chat as the LLM.
@@ B.2 LLM Prompts
Sadness: Sadness is a feeling of sorrow, unhappiness, or despondency.
Guilt: Guilt is a self-directed emotion that arises from a sense of wrongdoing or moral transgression.
@@
Fear: Fear is an emotion triggered by a perceived threat or danger.
Disgust: Disgust is an aversive emotion linked to feelings of revulsion, repulsion, or strong distaste. It arises in response to things that are offensive or unpleasant.
@@ B.3 Experimental Configurations
The detailed hyperparameter settings for online cascade learning are listed in Table 3 and 4. We tuned the hyperparameters using a grid search method on a separate validation set, which is a standard practice to avoid overfitting. In particular, we used the training set prepared for the knowledge distillation (as mentioned in Section 4.2) as our online cascade learning method’s validation set.
Specifically, for the hyperparameters $\beta$ and $\mu$, we observed that our experimental results are notably robust to variations in $\beta$. Regarding $\mu$, we tuned it specifically in the context of different cost budgets, which was essential for plotting Figure 3 and Figure 4 in our paper. By adjusting $\mu$, we were able to effectively manage the cost budgets to evaluate our method’s performance on cost-accuracy trade-offs, which is a critical aspect of our research objective.
Note that the “learning rate” in the table refers to the learning rates of the MLPs (in Section 3.2), not the models’. We used a consistent configuration for both online cascade learning and the distillation baselines by setting BERT-base’s batch size to 8, the learning rate to 0.00001, and the number of epochs to 5.
Table 3: Hyperparameter settings for online cascade learning experiments with GPT-3.5 Turbo as the LLM.
@@
Appendix C Discussion
Length Count Average Length GPT-3.5 Turbo Accuracy
52-664 4975 481.92 95.54%
664-843 5018 745.86 95.08%
843-1160 5003 985.05 93.96%
1160-1852 5001 1453.49 93.74%
1852-13704 5002 2953.95 92.44%
Total 25000 1325.07 94.15%
Table 5: GPT-3.5 Turbo’s classification accuracy across different IMDB review lengths. Longer inputs are typically more complex and thus have lower average accuracies.
C.1 Analysis of Training & Inference Cost Equilibrium
We conduct a thorough analysis of the training overhead for the models involved in our cascade to quantify the computational costs incurred in our approach. Below are the computation costs for training or inference over one sample (the computational costs of the confidence calibration MLP in Section 3.2, inference: $897\ Flops$, training: $1794\ Flops$, are negligible):
  • Logistic Regression training: $33.8\times 10^{4}Flops$ (floating-point operations).
  • Logistic Regression inference: $16.9\times 10^{4}Flops$.
  • BERT-base training: $18.5\times 10^{7}Flops$.
  • BERT-base inference: $9.2\times 10^{7}Flops$.
  • BERT-large training: $55.5\times 10^{7}Flops$.
  • BERT-large inference: $27.7\times 10^{7}Flops$.
[…]
Therefore, the computational costs regarding the deferral policy’s inference and training are minimal when comparing against the enormous LLM inference costs. In the real world, as the smaller models’ capabilities grow over time and the need for training decreases, the incurred model training overhead is also negligible compared to the tremendous LLM inference costs saved with our approach.
More formally, we can formulate the theoretical cost equilibrium as follows:
$100\%\cdot C=x\cdot M+(1-x)\cdot(M+2M+C),$
where the LHS refers to the inference overheads for using the LLM to process all queries, and the RHS comprises the maximum inference costs forusing the small models to handle $x\%$ of the queries (which are not deferred to the LLM), and the LLM inference costs, plus the inference & training costs for updating the small models when handling the rest $(1-x\%)$ queries. This equation can be further simplified to
$M=\frac{xC}{3-2x},$
where $C$ represents the LLM inference cost, $x$ is the proportion of queries handled by small models, and $M$ indicates small models’ aggregated inference costs.
For example, assuming $C=39.86\times 10^{15}Flops$, $x=0.5$, then $M\approx 9.95\times 10^{15}Flops$,typically refers to a total number of parameters around 17.5 Billion. That means, when using Llama 2 70B as the LLM, even if the smaller models can only handle 50% of the input queries, as long as the smaller models’ total number of parameters does not exceed 17.5B, our approach can still save costs.
C.2 Potential Overfitting in Larger Cascades
When scaling up the cascade, the task complexity may influence the appropriate cascade size, as suggested by the performance dynamics observed in Section 5.3. There are several factors that may affect this:
Task Complexity vs. Cascade Size
The HateSpeech dataset, despite its class imbalance, represents arelatively simple binary classification task where hate speech is often identifiable through specific keywords. Our findings, illustrated in Figure 6, show that a basic cascade of logistic regression and BERT-base can already handle  90% of the queries effectively, matching the performance of more complex models like GPT3.5 Turbo. This indicates that for simpler tasks, adding more models to the cascade might introduce unnecessary noise, complicating the deferral policy learning and thus degrading overall performance.
Capability Gap in Models
The performance impact of adding more models to a cascade also depends onthe capability gap between these models. In our larger cascade setup, we incorporated BERT-large alongside BERT-base. However, both models exhibit similar performance on the HateSpeech dataset, meaning the addition of BERT-large brought minimal benefit. This redundancy shows that the effectiveness of a cascade does not solely rely on adding more models, but rather on ensuring that each model contributes unique capabilities to the given task.
Improved Performance on More Complex Dataset
In contrast, for the more complex ISEAR dataset, the larger cascade that includes BERT-large outperformed the smaller cascade, despite the ISEAR dataset’s smaller size (3833) compared to HateSpeech (5352). This outcome supports the idea that as task complexity increases, a larger cascade can still better balance the cost-performance trade-off, leveraging the distinct strengths of each model in the cascade.
C.3 Challenges in Scaling Up Cascade
Scaling up the cascade for more complex task processing is technically feasible and compatible with our proposed framework, but can present notable challenges:
Confidence Measurement in Generative Tasks
For complex generative tasks, such as conversation, the measurement of confidence scores is also challenging. For example, traditional confidence estimation methods may not be well-suited for measuring LLMs’ token-by-token generation (Gupta et al., 2024). This difficulty in accurately assessing model confidence complicates the deferral policy learning within cascades, potentially leading to suboptimal routing of queries and affecting overall system efficiency.
Constraints with API-Based LLMs
More practically, many current LLMs, especially those accessible only via APIs, do not support fine-tuning. This restriction hinders our ability to tailor each LLM in the cascade to specific tasks in an online learning setting, presenting a significant limitation for customizing LLM-only cascades.
Computational Costs
The most important hurdle in a more complex cascade is the significant computational expense associated with online learning. For example, in an LLM-only cascade, updating LLMs in real-time (even if using efficient training techniques like LoRA) demands substantial resources, making it impractical for many applications. This computational burden not only affects scalability but also limits the frequency and extent to which LLMs can be updated, impacting the cascade’s adaptability and performance. As analyzed in Appendix C.1’s cost equilibrium, when the learnable model space grows too large, the training costs would offset the saved inference costs, which is against our initial motivation of enhancing cost-efficiency.